alones (alones):

zipdrix may you help here :3

alones (alones):

3x-y=12. What is the value of 8^x/2^y?

alones (alones):

Oh geez i spelled it wrong xP @zepdrix oh it's e there

OpenStudy (isaidavila):

find values for x and y I found that values X=7 and Y=9

OpenStudy (isaidavila):

8^7/2^9

OpenStudy (isaidavila):

(2^3)^7/2^9 2^21/2^9 2^12

Nnesha (nnesha):

rewrite 8^x in terms of base 2 and then apply the exponent rule \[\large\rm \frac{ x^m }{ x^n }=x^{m-n}\]

alones (alones):

Oh okay

Nnesha (nnesha):

so how would write 8^x in terms of base 2 ?

Nnesha (nnesha):

you*

alones (alones):

so 8/2?